OpenStudy (anonymous):

simplify (squrt(x)-7)/(x-49)

OpenStudy (anonymous):

= (sqrt(x)-7)/ (sqr(x)-7)((sqr(x)+7) = 1/((sqr(x)+7)

OpenStudy (anonymous):

= (sqrt(x) -7)/(x-49)

OpenStudy (anonymous):

that's possible : \[(\sqrt{x}-7)/(\sqrt{x}-7)(\sqrt{x}+7) = 1/(\sqrt{x}+7)\]

OpenStudy (anonymous):

also you have as: \[[1/(\sqrt{x}+7) ]*[ (\sqrt{x}-7)/(\sqrt{x}-7)] = (\sqrt{x}-7)/(x-49)\]